Grade 12th passAlgebra

If you subtract ½ from a number and multiply the result by ½, you get 1/8 what is the number?

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let the number be x.
According to the question,
(x–1/2)× 1⁄2 = 1/8
x/2–1⁄4=1/8
x/2=1/8+1⁄4
x/2=1/8+2/8
x/2 = (1+ 2)/8
x/2 = 3/8
x = (3/8) × 2
x = 3⁄4

Thanks